OpenStudy (anonymous):

Medal and fanning !!!!!!!!! thank you ^_^ 1. x2 – 8x + 15 2. a2 – a – 20 3. a2 – 5a – 20 4. a2 + 12ab + 27b2 5. 2a2 + 30a + 100

OpenStudy (anonymous):

1. x^2 - 8x + 15 x^2 - 5x - 3x + 15 x(x - 5) -3(x - 5) (x - 3)(x - 5) 2. a^2 - a - 20 a^2 - 5a + 4a - 20 a(a - 5) 4(a - 5) (a + 4)(a - 5) 3. a^2 - 5a - 20 Unfactorable. 4. a^2 + 12ab + 27b^2 a^2 + 9ab + 3ab + 27b^2 a(a + 9b) 3b(a + 9b) (a + 3b)(a + 9b) 5. 2a^2 + 30a + 100 2a^2 + 20a + 10a + 100 2a(a + 10) 10(a + 10) (2a + 10)(a + 10)

OpenStudy (anonymous):

@Hero

hero (hero):

From where did you learn this factoring method? It's the exact same kind I use.

hero (hero):

By the way, a^2 - 5a - 20 is factorable

OpenStudy (anonymous):

Flvs but i think two of them are wrong

OpenStudy (anonymous):

Ohh can you show me how hero? thank you

hero (hero):

For Number 5, the factoring steps are this: 2a^2 + 30a + 100 = 2(a^2 + 15a + 50) = 2(a^2 + 10a + 5a + 50) = 2((a(a + 10) + 5(a + 10)) = 2((a + 10)(a + 5)) = 2(a + 10)(a + 5)

OpenStudy (anonymous):

Thanks thats what i thought too but please tell me which ones are wrong and which ones are right and if whatever ones i got wrong please show me the right way thanks hero

hero (hero):

Technically you didn't get any of the wrong. I was just saying that number 3 is factorable.

OpenStudy (anonymous):

Oh ok thanks but can you explain how its factorable because I don't know how

OpenStudy (anonymous):

and do you think for number two the sign needs to be changed?

hero (hero):

The sign for number 2 doesn't need to be changed because \(-5 \times 4 = -20\) \(-5 + 4 = -1\) So you found the right two numbers

hero (hero):

Number 3 would take a bit more effort to explain.

OpenStudy (anonymous):

Ok thank you but could you please help for number three

hero (hero):

a^2 - 5a - 20 You need two numbers \(m\) and \(n\) that multiply to get -20 yet add to get -5 In other words \(mn = -20\) \(m + n = -5\) It's a bit much to show here.

OpenStudy (anonymous):

hows about 4*-5 equals =-20

hero (hero):

-5 + 4 doesn't add up to -5

OpenStudy (anonymous):

and 15-20=-5

OpenStudy (anonymous):

Ohhh

hero (hero):

m and n have to be the same

OpenStudy (anonymous):

Then i do not think its possible

hero (hero):

It's possible

OpenStudy (anonymous):

Hmm

hero (hero):

What I mean is m and n have to be consistent for both equations.

hero (hero):

Like what I showed you above for number two.

OpenStudy (anonymous):

Yeah i see but i don't know which numbers would work

hero (hero):

Trust me, it's not worth knowing at your level right now.

OpenStudy (anonymous):

Loll I just searched that one up every ones saying its prime

OpenStudy (anonymous):

Some one says this It is prime and will not factor. If you are looking for roots (zeros), then you can complete the square or use the quadratic formula. Let's complete the square. a^2 - 5a - 20 = 0 Add 20 to both sides: a^2 - 5a = 20 Add (-5/2)^2 to both sides to complete the square: a^2 - 5a + 25/4 = 20 + 25/4 Factor the left side, and simplify the right: (a - 5/2)^2 = 105/4 Take the square root of both sides, but be careful not to lose the negative root: a - 5/2 = (±√105) / 2 Add 5/2 to both sides: a = (5±√105)/2

OpenStudy (anonymous):

what do you think about it @Hero

hero (hero):

Of course you can find the values that way. But I was referring to the factoring method.

hero (hero):

I'll show it to you later. I'm going to go take a nap.

OpenStudy (anonymous):

Lol Ok have a good nap

OpenStudy (anonymous):

@SOS101